# Analysis of ion chain sympathetic cooling and gate dynamics

### v1

@@ I Introduction
Trapped-ion quantum computers are one near-term implementation of quantum computing, where ions trapped in electric fields form chains along the electric potential’s weakest axis Häffner et al. (2008); Bruzewicz et al. (2019). A significant challenge to high-fidelity entangling gate operations in trapped ions is motional heating that causes significant gate errors and decoherence in long chains of ions Cetina et al. (2022a); Sutherland et al. (2022); Semenin et al. (2022). One proposed method to combat this heating is with sympathetic cooling Itano et al. (1991); Kielpinski et al. (2000); Mao et al. (2021); Aikyo et al. (2023). Introducing ancillary coolant ions into the chain interspersed between the qubit ions, the Coulomb coupling between the ions can be used to cool the qubits indirectly by laser-cooling the coolants. Thus, the electronic state of the qubits is not destroyed while the chain is being cooled.
However, this strategy comes with trade-offs; firstly, adding too many coolant ions can increase the length of the chain, making it more susceptible to thermal noise due to the reduction in motional mode frequency Cetina et al. (2022b); Brownnutt et al. (2015). In addition, gates cannot be run while ions are being cooled due to the use of ion motion to execute the gate; thus, if the chain is cooled for too long, other errors can dominate the noise budget, such as dephasing effects.
@@ II Modeling chain dynamics
We define a dimensionless potential $V(\vec{u})$ normalized to MHz frequency units Egan (2021), i.e. $V=U/E_{0}$ and $\vec{u}=\vec{x}/d_{0}$, where
@@
Ions are often trapped in an equally-spaced chain, allowing a single optical device with equally spaced channels such as an acousto-optic modulator or array of fibers to optically address all ions easily Egan et al. (2021); Binai-Motlagh et al. (2023). To create an equispaced chain, a quadratic potential is insufficient. We compensate by adding in a quartic term to our potential. Thus, in the direction of the chain, ions can be trapped at normalized positions $u_{i}$ in a symmetric quartic potential of the form:
@@ II.1 Semiclassical model
At equilibrium, the ions are cold enough that the potential can be linearized, allowing the dynamics of the ions in the chain to be well-approximated as a set of coupled oscillators Egan (2021); Guggemos et al. (2015). We use Eq. 1 to generate an equation of motion for this coupled oscillator system by taking the Hessian $K$ of $V(\mathbf{u})$, a set of effective spring constants and coupling coefficients between the ions. The solutions to such a system are well-known and form the normal modes of this chain. The lowest-frequency of these modes in the axial direction is the center-of-mass (COM) mode. In this mode, all the ions move in phase with each other. The COM mode is generally the largest contributor to axial heating in ion chains, as low-frequency electric field noise can be approximated as a DC electric field displacing the entire chain in the same direction.
We can model cooling by extending the semiclassical model to include a damping force Metcalf and van der Straten (2003); Guggemos et al. (2015); Morigi and Walther (2001). The attenuation of a particular mode scales with the mode amplitude, yielding exponential decay of the mode amplitude over time. The damping matrix, $\Gamma$, represents the damping force imparted upon the ions by the cooling light.
@@
By using an ansatz $\mathbf{x}(t)=\mathbf{v}e^{i\omega t}$, we can derive the eigenvalues of $K$$-\omega_{i}^{2}$} and its corresponding eigenvectors $\mathbf{v}_{i}$}. These eigenvectors represent the modes of this oscillator system, while the eigenvalues represent their frequencies. These dynamics can be seen in figure 2, with linear increase in mode amplitude when heating and exponentially decay of the mode when cooling.
Figure 2: Cooling dynamics. (a) COM mode amplitude, and (b) mean XX gate fidelity with a $60\%$ duty cycle (60% of each cycle is spent cooling the chain) over three cooling cycles. Shaded sections correspond to sympathetic cooling intervals, while unshaded sections represent intervals where gates are being run. Parameters used in this simulation include a chain length of 15 ions with $4.4~{}\mu$m spacing and 640 kHz cooling Rabi frequency.
@@ II.2 Coolant placement - perturbative approach
To find the eigenvalues of this system, we can use a perturbative approach for two reasons. Firstly, the damping force present on single ions is small relative to the mode frequencies. The timescale for COM mode cooling is on the order of milliseconds, which corresponds to a cooling rate of  $10^{2}$ Hz, in comparison to the mode frequency of  $10^{5}$ Hz Nop et al. (2021); Egan (2021). Secondly, $\Gamma$ is a sparse, diagonal matrix, with only $N_{C}$ nonzero entries for $N_{C}$ coolants in the chain, each at indices $C$. As such, the eigenvalues and eigenvectors of the damped system will be close enough to those of the unperturbed system, and thus we can approximate the new eigenvectors and eigenvalues based on the original modes and frequencies of the system.
Because the cooling rate is constant for all ions, we can rewrite $\Gamma=\gamma P$, where $P$ is some projector matrix with ones at specific coolant positions along the diagonal. Next, we use a similar ansatz to find our eigenvalues and eigenvectors. By using $\mathbf{x}(t)=\mathbf{w}e^{zt}$, we will get eigenvectors $\mathbf{w}_{i}$ and complex eigenvalues $z_{i}^{2}$. $\Im\{z_{i}\}$ corresponds to the oscillation frequency of the $i$th damped mode, while $\Re\{z_{i}\}$ corresponds to the cooling rate of the $i$th damped mode. Because we can solve this system perturbatively with respect to $\gamma$, we set $\mathbf{w}_{i}=\sum_{k=0}^{\infty}\gamma^{k}\mathbf{w}_{i}^{(k)}$, and $z_{i}^{2}=s_{i}=\sum_{k=0}^{\infty}\gamma^{k}s_{i}^{(k)}$. In addition, $\mathbf{w}_{i}^{(0)}=\mathbf{v}_{i}$, while $s_{i}^{(0)}=-\omega_{i}^{2}$. We substitute the ansatz into Eqn 2:
@@
To check the validity of this solution, we compare it to the computed eigenvalues of the damped system given in equation 2. The perturbative approximation was accurate to the computed eigenvalues within $10^{-8}\gamma$.
@@ II.3 Cooling limit
To compute the cooling limit for a particular mode, we use our previous approximation for the COM mode cooling rate in conjunction with results from Cetina et al. (2022a), which show empirically that the mode heating rate $h$ is related to axial chain frequency by a power law of the form
$h=\frac{dn}{dt}=A\omega_{0}^{-1-\alpha}+B$ (6)
Combining these results with Eq. 3, we get:
@@
$=\frac{A\omega_{0}^{-1-\alpha}+B}{\frac{\omega}{\sqrt{2}}\sqrt{\sqrt{1+g_{0}^{2}\left(\sum_{k\in P}|v_{0k}|^{2}\right)^{2}}-1}}$ (7)
$\approx\frac{A\omega_{0}^{-2-\alpha}}{\frac{1}{\sqrt{2}}\sqrt{\sqrt{1+g_{0}^{2}\left(\sum_{k\in P}|v_{0k}|^{2}\right)^{2}}-1}}$ (8)
$\approx\frac{2A\omega_{0}^{-2-\alpha}}{g_{0}\left(\sum_{k\in P}|v_{0k}|^{2}\right)}$ (9)
One observation from the above expression is that our cooling performance becomes much better as we increase our COM mode frequency, as this improves our heating rate as well as our cooling rate. The heating rate decreases due to its scaling with $\omega^{-1-\alpha}$, while the cooling rate increases approximately linearly in $\omega$. On the other hand, increasing the individual ion damping parameter will increase our cooling capabilities linearly, while not affecting our heating rate at all.
@@
Although we don’t have an analytical expression for $\omega_{0}$ or $v_{0k}$ in a quartic potential, the highest mode participation factors for the COM mode in a quartic potential come from the ions at the center of the chain (see Figure 1), we know that the optimal ion positions must be at the center of the chain.
Finally, we observe that many of the quantities in the above expression can be determined analytically for a quadratic potential. For a normalized quadratic potential given by $V(u_{i})=X_{2}u_{i}^{2}$, our normalization gives $\omega_{0}=(2\pi\times 1~{}\text{MHz})\sqrt{X_{2}}$, while $v_{0k}=N^{-1/2}$ for all positions in an $N$-ion chain Morigi and Walther (2001). The latter reduces the sum $\sum_{k\in P}|v_{ik}|^{2}$ to $\frac{N_{C}}{N}$. This provides an analytical cooling limit for any configuration of an ion chain trapped in a quadratic potential.
However, this also provides us with an analytical upper-bound for a quartic potential. In the limit where a few coolants are placed optimally at the center of a chain trapped in a quartic potential, the higher mode participation factors at the center of the chain ensure that the sum $\sum_{k\in P}|v_{ik}|^{2}>\frac{N_{C}}{N}$, while the inclusion of a nonzero $X_{4}$ term will only make the potential tighter, thus causing a higher COM frequency. As such, because we can show that the analytical quadratic heating rate is lower than its quartic counterpart and the analytical quadratic cooling rate is lower than than the quartic cooling rate, the quadratic cooling rate can act as a soft upper bound on the cooling limit of a corresponding quartic chain.
@@ III Simulation and Analysis
Now we apply the above theoretical model to find optimal cooling parameters for long-chain trapped-ion quantum computers through computational simulation by maximizing mean gate fidelity over time. We first compute the eigenfrequencies and eigenmodes of the system given in equation 2 in both the undamped ($\Gamma=O$) and damped ($\Gamma=\gamma P$) regimes for a particular cooling configuration. We then use the eigenvalues of equation 2 to compute the cooling rate of the COM mode, while using equation 6 to compute the COM heating rate. These rates are subsequently used to simulate COM mode dynamics, as shown in Figure 2(a). Finally, we use results found in Cetina et al. (2022b) to compute XX gate fidelity from the resultant mode dynamics, seen in Figure 2(b).
@@
We examine two separate sets of parameters to optimize: chain parameters such as the number of coolants, and cooling parameters such as cooling duty cycle.
As a proof of concept, the simulation was used to find optimal coolant ion configurations and cooling schedules for a quantum algorithm given in Wang et al. (2021) run on the trapped-ion quantum computer specified in Cetina et al. (2022a).
@@ III.1 Parameters
The parameters for the initial simulation were largely derived from Cetina et al. (2022a). Heating rates for the ion chain were derived using the relation in Eq. 6 with $\alpha=0.8$, $A\approx 8.2\times 10^{17}\text{ s}^{-1}\times(1\text{ rad/s})^{2+\alpha}$, and $B=0.9\text{ s}^{-1}$, as these parameters yielded good agreement with experimental data. We conduct simulations using cooling Rabi frequencies of 180, 275, and 640 kHz.
@@ III.2 Chain optimization
The impact of COM mode frequency in a fixed potential (and, by proxy, chain length) has been examined at length in Cetina et al. (2022a). Thus, in this study, we examine cooling performance through the lens of mean XX gate fidelity and how it is impacted by the number of coolants and chain length. Because these parameters are determined by the physical characteristics of the trapped-ion system, these results are largely dependent on the number of ions used to run a circuit, and are independent of circuit length and gate type.
We first compute the mean gate fidelity of a circuit run on the chain as we increase the number of coolants in the chain with a fixed number of qubits. As seen in Fig. 3(a), the optimal number of coolants depends on the cooling duty cycle. Having fewer than the optimal number of coolants in the chain causes the coolants to not couple strongly enough to the COM mode, leading to insufficient cooling of the mode, lowering gate fidelity. However, more than 6 coolants causes the COM mode frequency to decrease due to chain length (as shown in Cetina et al. (2022a)), and thus increases the overall heating rate of the COM mode, making the chain harder to cool. As the cooling time for each cooling cycle is increased, the optimal number of coolants varies. Thus, the optimal chain configuration is largely dependent on cooling parameters, and may be determined computationally for fixed cooling cycle parameters. We also examine a more general case for a chain of fixed length in Appendix A. We find trivially that the cooling limit asymptotically approaches a fixed value, achieved when the chain is entirely composed of 172Yb+ coolants, as expected.

### v2

@@ I Introduction
Trapped-ion quantum computers are one near-term implementation of quantum computing, where ions trapped in electric fields form chains along the electric potential’s weakest axis [1, 2]. A significant challenge to high-fidelity entangling gate operations in trapped ions is motional heating that causes significant gate errors and decoherence in long chains of ions [3, 4, 5]. One proposed method to combat this heating is with sympathetic cooling [6, 7, 8, 9]. Introducing ancillary coolant ions into the chain interspersed between the qubit ions, the Coulomb coupling between the ions can be used to cool the qubits indirectly by laser-cooling the coolants. Thus, the electronic state of the qubits is not destroyed while the chain is being cooled.
However, this strategy comes with trade-offs; firstly, adding too many coolant ions can increase the length of the chain, making it more susceptible to thermal noise due to the reduction in motional mode frequency [10, 11]. In addition, gates cannot be run while ions are being cooled due to the use of ion motion to execute the gate; thus, if the chain is cooled for too long, other errors can dominate the noise budget, such as dephasing effects.
@@ II Modeling chain dynamics
We define a dimensionless potential $V(\vec{u})$ normalized to MHz frequency units [12], i.e. $V=U/E_{0}$ and $\vec{u}=\vec{x}/d_{0}$, where
@@
Ions are often trapped in an equally-spaced chain, allowing a single optical device with equally spaced channels such as an acousto-optic modulator or array of fibers to optically address all ions easily [13, 14]. To create an equispaced chain, a quadratic potential is insufficient. We compensate by adding in a quartic term to our potential. Thus, in the direction of the chain, ions can be trapped at normalized positions $u_{i}$ in a symmetric quartic potential of the form:
@@ II.1 Semiclassical model
At equilibrium, the ions are cold enough that the potential can be linearized, allowing the dynamics of the ions in the chain to be well-approximated as a set of coupled oscillators [12, 15]. We use Eqn. 1 to generate an equation of motion for this coupled oscillator system by taking the Hessian $K$ of $V(\mathbf{u})$, a set of effective spring constants and coupling coefficients between the ions. The solutions to such a system are well-known and form the normal modes of this chain. The lowest-frequency of these modes in the axial direction is the center-of-mass (COM) mode. In this mode, all the ions move in phase with each other. The COM mode is generally the largest contributor to axial heating in ion chains, as low-frequency electric field noise can be approximated as a DC electric field displacing the entire chain in the same direction.
We can model cooling by extending the semiclassical model to include a damping force [16, 15, 17]. The attenuation of a particular mode scales with the mode amplitude, yielding exponential decay of the mode amplitude over time. The damping matrix, $\Gamma$, represents the damping force imparted upon the ions by the cooling light.
@@
By using an ansatz $\mathbf{x}(t)=\mathbf{v}e^{i\omega t}$, we can derive the eigenvalues of $K$$-\omega_{i}^{2}$} and its corresponding eigenvectors $\mathbf{v}_{i}$}. These eigenvectors represent the modes of this oscillator system, while the eigenvalues represent their frequencies. These dynamics can be seen in Fig. 2, with linear increase in mode amplitude when heating and exponential decay of the mode when cooling.
Figure 2: Cooling dynamics. (a) simulated COM mode amplitude (with analytical cooling limit described in section II illustrated), and (b) simulated mean XX gate fidelity with a $60\%$ duty cycle (60% of each cycle is spent cooling the chain) over three cooling cycles. Shaded sections correspond to sympathetic cooling intervals, while unshaded sections represent intervals where gates are being run. Parameters used in this simulation include a chain length of 23 ions with $3.7~{}\mu$m spacing and 640 kHz cooling Rabi frequency.
@@ II.2 Coolant placement - perturbative approach
To find the eigenvalues of this system, we can use a perturbative approach for two reasons. Firstly, the damping force present on single ions is small relative to the mode frequencies. The timescale for COM mode cooling is on the order of milliseconds, which corresponds to a cooling rate of  $10^{2}$ Hz, in comparison to the mode frequency of  $10^{5}$ Hz [18, 12]. Secondly, $\Gamma$ is a sparse, diagonal matrix, with only $N_{C}$ nonzero entries for $N_{C}$ coolants in the chain, each at indices $C$. As such, the eigenvalues and eigenvectors of the damped system will be close enough to those of the unperturbed system, and thus we can approximate the new eigenvectors and eigenvalues based on the original modes and frequencies of the system.
Because the cooling rate is constant for all coolant ions, we can rewrite $\Gamma=\gamma P$, where $P$ is some projector matrix with ones at specific coolant positions along the diagonal. Next, we use a similar ansatz to find our eigenvalues and eigenvectors. By using $\mathbf{x}(t)=\mathbf{w}e^{zt}$, we will get eigenvectors $\mathbf{w}_{i}$ and complex eigenvalues $z_{i}^{2}$. $\Im\{z_{i}\}$ corresponds to the oscillation frequency of the $i$th damped mode, while $\Re\{z_{i}\}$ corresponds to the cooling rate of the $i$th damped mode. Because we can solve this system perturbatively with respect to $\gamma$, we set $\mathbf{w}_{i}=\sum_{k=0}^{\infty}\gamma^{k}\mathbf{w}_{i}^{(k)}$, and $z_{i}^{2}=s_{i}=\sum_{k=0}^{\infty}\gamma^{k}s_{i}^{(k)}$. In addition, $\mathbf{w}_{i}^{(0)}=\mathbf{v}_{i}$, while $s_{i}^{(0)}=-\omega_{i}^{2}$. We substitute the ansatz into Eqn. 2:
@@
To check the validity of this solution, we compare it to the computed eigenvalues of the damped system given in Eqn. 2. The perturbative approximation was accurate to the computed eigenvalues within $\approx 10^{-6}\gamma$ for values of $\gamma$ corresponding to the range of cooling Rabi frequencies employed for Doppler cooling of trapped ions.
@@ II.3 Cooling limit
To compute the cooling limit for a particular mode, we use our previous approximation for the COM mode cooling rate in conjunction with results from [3], which show empirically that the mode heating rate $h$ is related to axial mode frequency by a power law of the form
$h=\frac{dn}{dt}$ $=D\omega_{0}(A_{0}\omega_{0}^{-2-\alpha}+B_{0})$ (6)
$=(A\omega_{0}^{-1-\alpha}+B\omega_{0})$ (7)
In Eqn. 6, $D$ is a normalization factor derived from [3] and is dependent on chain parameters such as trapping potential and cooling Rabi frequency. Thus, in this work, we fix $A_{0}$ and $B_{0}$ as chain-independent parameters when modeling heating dynamics.In long chains, COM frequencies are generally on the order of 105 Hz [3]. As such, so long as the axial COM frequency remains in this regime, the contribution of the $B\omega_{0}$ term is relatively small.
Combining these results with Eqn. 3, we get:
@@
$=\frac{A\omega_{0}^{-1-\alpha}+B\omega_{0}}{\frac{\omega}{\sqrt{2}}\sqrt{\sqrt{1+g_{0}^{2}\left(\sum_{k\in C}|v_{0k}|^{2}\right)^{2}}-1}}$ (8)
$\approx\frac{A\omega_{0}^{-2-\alpha}+B}{\frac{1}{\sqrt{2}}\sqrt{\sqrt{1+g_{0}^{2}\left(\sum_{k\in C}|v_{0k}|^{2}\right)^{2}}-1}}$ (9)
$\approx\frac{2(A\omega_{0}^{-2-\alpha}+B)}{g_{0}\left(\sum_{k\in C}|v_{0k}|^{2}\right)}$ (10)
$=\frac{2(A\omega_{0}^{-1-\alpha}+B\omega_{0})}{\gamma\left(\sum_{k\in C}|v_{0k}|^{2}\right)}$ (11)
One observation from the above expression is that our cooling performance becomes much better as we increase our COM mode frequency, as this decreases our heating rate while increasing our cooling rate. In contrast, increasing the individual ion damping parameter will increase our cooling capabilities linearly, while not affecting our heating rate at all.
@@
Figure 3: Computational verification of theoretical observations. (a) Cooling limit with respect to coolant proportion and COM frequency. The increase in cooling limit in the vertical direction corresponds with an increase in heating rate due to a lower COM frequency, while the decrease in cooling limit in the rightward direction corresponds to a larger number of coolants in the chain. The cooling limit plateaus as the chain fills with coolants. (b) Simulated COM mode cooling rates for all possible configurations of two coolants in a 15-ion chain. Optimal placement is found with ions placed at positions (0, 1) or (-1, 0), in good agreement with our theoretical observation that optimal cooling performance occurs with ions placed at the center of the chain.
Although we don’t have an analytical expression for $\omega_{0}$ or $v_{0k}$ in a quartic potential, the highest mode participation factors for the COM mode in a quartic potential come from the ions at the center of the chain (see Fig. 1), and thus we know that the optimal ion positions must be at the center of the chain.
We also use computational methods to verify the accuracy of our analytical observations by examining the correspondence between the perturbative definitions of the cooling and the computationally simulated cooling limit. Fig. 2(a) shows the accuracy of the cooling limit. Our analytical results show good agreement to the simulated dynamics, accurately estimating the limit of a particular cooling scheme given its geometry. In addition, we also found that the perturbative approximation of the cooling rate only had a relative error of $<2.5\times 10^{-4}$ for all $\gamma<10^{-3}$.
We computationally verify our observation that optimal coolant placement is at the center of a chain by simulating every possible configuration of two coolant ions in a 15-ion chain, seen in Fig. 3(b). Cooling performance was maximized with centrally-placed coolants (i.e. at positions (-1, 0) or (0, 1)), yielding a cooling limit of 29 quanta. This is in good agreement with our theoretical observations, as configurations that perform best (i.e. minimize the cooling limit) are centrally placed, while configurations that perform worse place coolant ions towards the edge of the chain.
However, this perturbative approximation will likely not hold at higher values of $\gamma$ due to hybridization of the undamped COM mode into other higher-order modes of damped motion that arise in long ion chains. In the small $\gamma$ regime, the perturbation does not significantly change the modes of the damped system when compared to those of the undamped system. As such, each undamped mode largely couples to its damped counterpart, including the undamped COM mode.
Finally, we observe that many of the quantities in the above expression can be determined analytically for a quadratic potential. For a normalized quadratic potential given by $V(u_{i})=X_{2}u_{i}^{2}$, our normalization gives $\omega_{0}=(2\pi\times 1~{}\text{MHz})\sqrt{X_{2}}$, while $v_{0k}=N^{-1/2}$ for all positions in an $N$-ion chain [17]. The latter reduces the sum $\sum_{k\in P}|v_{ik}|^{2}$ to $\frac{N_{C}}{N}$. This provides an analytical cooling limit for any configuration of an ion chain trapped in a quadratic potential.
However, this also provides us with an analytical upper-bound for a quartic potential. In the limit where a few coolants are placed optimally at the center of a chain trapped in a quartic potential, the higher mode participation factors at the center of the chain ensure that the sum $\sum_{k\in P}|v_{ik}|^{2}>\frac{N_{C}}{N}$, while the inclusion of a nonzero $X_{4}$ term will only make the potential tighter, thus causing a higher COM frequency. As such, because we can show that the analytical quadratic heating rate is higher than its quartic counterpart and the analytical quadratic cooling rate is lower than the quartic cooling rate, the quadratic cooling rate can act as a soft upper bound on the cooling limit of a corresponding quartic chain.
@@ III Simulation and Analysis
Now we apply the above theoretical model to find optimal cooling parameters for long-chain trapped-ion quantum computers through computational simulation by maximizing mean gate fidelity over time. We first compute the eigenfrequencies and eigenmodes of the system given in Eqn. 2 in both the undamped ($\Gamma=0$) and damped ($\Gamma=\gamma P$) regimes for a particular cooling configuration. We then use the eigenvalues of Eqn. 2 to compute the cooling rate of the COM mode, while using Eqn. 6 to compute the COM heating rate. These rates are subsequently used to simulate COM mode dynamics, as shown in Fig. 2(a). Finally, we use results found in [10] to compute XX gate fidelity from the resultant mode dynamics, seen in Fig. 2(b).
@@
Gate fidelity depends on the shape of the individual addressing beams used in the specific system, in addition to the COM mode frequency. For these properties, we utilize parameters used in situ in [3]. We examine two separate sets of parameters to optimize: chain parameters such as the number of coolants, and cooling parameters such as cooling duty cycle.
As a proof of concept, the simulation was used to find optimal coolant ion configurations and cooling schedules for a quantum algorithm given in [20] run on the trapped-ion quantum computer specified in [3].
@@ III.1 Parameters
The parameters for the initial simulation were largely derived from [3]. Heating rates for the ion chain were derived using the relation in Eqn. 6 with $\alpha=0.8$, $A_{0}\approx 8.2\times 10^{17}\text{ s}^{-1}\times(1\text{ rad/s})^{2+\alpha}$, and $B_{0}=0.9\text{ s}^{-1}$, as these parameters yielded good agreement with experimental data. We conduct simulations using cooling Rabi frequencies of 180, 275, and 640 kHz. The corresponding cooling rates were also determined to be $\gamma$ = $1.387\times 10^{-5}$, $3.468\times 10^{-5}$, and $5.328\times 10^{-5}$ respectively, as these parameters also gave good agreement with the cooling dynamics described in [3].
@@ III.2 Chain optimization
The impact of COM mode frequency in a fixed potential (and, by proxy, chain length) has been examined at length in [3]. Thus, in this study, we examine cooling performance through the lens of mean XX gate fidelity and how it is impacted by the number of coolants and chain length. Because these parameters are determined by the physical characteristics of the trapped-ion system, these results are largely dependent on the number of ions used to run a circuit, and are independent of circuit length and gate type.
We first compute the mean gate fidelity of a circuit run on the chain as we increase the number of coolants in the chain with a fixed number of qubits. As seen in Fig. 4(a), the optimal number of coolants depends on the cooling duty cycle. Having fewer than the optimal number of coolants in the chain causes the coolants to not couple strongly enough to the COM mode, leading to insufficient cooling of the mode, lowering gate fidelity. However, more than 6 coolants causes the COM mode frequency to decrease due to chain length (as shown in [3]), and thus increases the overall heating rate of the COM mode, making the chain harder to cool. As the cooling time for each cooling cycle is increased, the optimal number of coolants varies. Thus, the optimal chain configuration is largely dependent on cooling parameters, and may be determined computationally for fixed cooling cycle parameters. For the purposes of this study, it’s sufficient to assume a fixed electric potential for ion trapping. However, below, we also briefly examine how changing aspects of the trap potential (in particular, the COM mode frequency) impacts cooling performance.
Rather than adding coolants to the chain for a fixed number of qubits, we examine the ratio of coolants to total ions in the chain. In so doing, we examine the effect of increasing the cooling rate of the mode by maximizing the total mode participation over all coolants (i.e., maximizing $\sum_{k\in P}|v_{0k}|^{2}$ in Eqn. 3). We test 21-ion chains with differing COM frequencies as we add coolants starting at the center of the chain, subsequently assessing their cooling performance. Fig. 3(a) shows the corresponding cooling limit of a chain for a fixed COM frequency and coolant filling fraction. The theoretical prediction for our cooling limit holds well here, with the cooling limit decreasing as the COM frequency increases. In addition, the cooling performance plateaus as the coolant filling fraction approaches 1. This is indicative of the trends observed in Fig. 4, in which cooling performance plateaus and eventually decreases as chain length increases.